A voltmeter connected in series with a resistance R1 to a circuit indicates a voltage V1=198V. When a series resistor R2=2R! is used,the voltmeter indicates a voltage V2=180V.If the resistance of the voltmeter is Rv=900 ohm then,the applied voltage across A and B is

Grade:12

1 Answers

rajula vineet reddy
9 years ago
the answer is 220v....let V be the applied voltage for the first case the eqn is 198=(Rv*V)/(R1+Rv) and fo rthe second case 180=(Rv*V)/(R2+Rv) solve the two u get the answer.